OpenStudy (anonymous):

An object of mass m=1.0 kg experiences a force of the following mathematical form- F(t)=ke^(-ct) where k=6.0 N, and c=0.101/s. At time t=0, the object has a velocity of v=2.0m/s and is at the origin. Where is this object after 20 seconds?

OpenStudy (fifciol):

Start from this equation: \[F=m \frac{ dv }{dt }\] Can you determine v out of it?

OpenStudy (fifciol):

\[Fdt=mdv \Rightarrow dv=\frac{ 1 }{ m }Fdt\] integrate both sides: \[\int\limits dv=\frac{ 1 }{m }\int\limits Fdt=\frac{ 1 }{ m }\int\limits ke^{-ct}dt\] so \[v=\frac{ k }{ m }\int\limits e^{-ct}dt\] can you solve this integral?

OpenStudy (fifciol):

@schi1 are you still there?

OpenStudy (anonymous):

where has the k-exp(ct) come from @schi1?

OpenStudy (anonymous):

umm the formula is- \[F(x)=ke ^{-ct}\] where e is the constant e

OpenStudy (anonymous):

and yes I did solve the integral, thank you @Fifciol

OpenStudy (fifciol):

what is the result then? Or you've already solved the problem?